OpenStudy (anonymous):

derivative y=(2)^(5x+3)

hartnn (hartnn):

can you derivate a^x ?

OpenStudy (anonymous):

I'm not sure.. I need help like step by step through it too btw.

hartnn (hartnn):

yes, do u know what is chain rule ?

OpenStudy (anonymous):

yes. A'(B)*B'

hartnn (hartnn):

thats correct. now, \(\huge \frac{d}{dx}a^x=a^x\ln a\) so what will be derivative of 2^(5x+3) =?

OpenStudy (anonymous):

2^(5x+3)*ln(2)

hartnn (hartnn):

let me give you 1st step, \(y=2^{5x+3} \\ \text{differentiate both sides w.r.t x} \\ y'=2^{5x+3}\ln2(5x+3)' \quad \text{using chain rule, got this ?}\)

hartnn (hartnn):

\(y'=2^{5x+3}(\ln2)(5x+3)'=? \)

OpenStudy (anonymous):

5ln(2)(2)^(5x+3)

hartnn (hartnn):

absolutely correct! thats your final answer....

hartnn (hartnn):

did u understand ?

OpenStudy (anonymous):

oh thanks! ;D

OpenStudy (anonymous):

yeah :) it makes more sense now!

hartnn (hartnn):

welcome ^_^